What is the pH of a saturated solution of silver hydroxide, given Ksp = 1.53×10-8?


1
Expert's answer
2021-10-14T04:03:59-0400

"AgOH\leftrightharpoons" Ag+ + OH-

KSp= [Ag+][OH-]

1.53×10-8=(s)(s)

1.53×10-8=s2

s=(1.53×10-8)1/2=1.23×10-4 M=[OH-]=[Ag+]

pH + pOH =14

pH = 14 - pOH = ?

First we calculate pOH= -log(OH-)=

pOH = -log(1.23×10-4)=3.91

pH + pOH =14

pH + 3.91=14

pH = 14-3.91=10.08 is the pH of a saturated solution of silver hydroxide